A line passing through (9,8) and (7,y) is parallel to a line with slope -1. What is the value of y?
Thank You!

Answer by Nate(3500)

Use point-slope form:
y - yp = m(x - xp) where the variable and 'p' are the points *use (9,8)*
y - 8 = (-1)(x - 9)
y - 8 = -x + 9
y = -x + 17
Now, plug in (7,y):
y = -x + 17
y = -7 + 17
y = 10
So, the ordered pair is: (7,10)